# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ Abstract
Gaussian mixture models (GMMs) are fundamental statistical tools for modeling heterogeneous data. Due to the nonconcavity of the likelihood function, the Expectation-Maximization (EM) algorithm is widely used for parameter estimation of each Gaussian component. Existing analyses of the EM algorithm’s convergence to the true parameter focus primarily on either the two-component case or multi-component settings with both known mixing probabilities and known, isotropic covariance matrices.
In this work, we establish the minimax optimal rate of convergence of the EM algorithm for multi-component GMMs in full generality. The required separation condition between Gaussian components for EM to converge is the weakest known to date. We develop two distinct analytical approaches, each tailored to a different regime of separation, reflecting two complementary perspectives on the use of EM: parameter estimation and clustering. As a byproduct of our analysis, we show that the EM algorithm, when used for community detection, also achieves the minimax optimal rate of misclustering error under milder separation conditions than spectral clustering and Lloyd’s algorithm, an interesting result in its own right. Our analysis allows the number of components, the minimal mixing probabilities, the separation between Gaussian components as well as the dimension to grow with the sample size. Simulation studies corroborate the theoretical findings.
Keywords: Gaussian mixture models, minimax optimal estimation, EM algorithm, community detection, clustering.
@@ 1 Introduction
Finite mixture models are fundamental statistical tools for modeling heterogeneous data [28, 32] and have found wide-ranging applications in fields such as biology, economics, image processing, and natural language processing, to name just a few. These models assume that the observed data arise from a mixture of several latent sub-populations, each governed by its own distribution [31, 39]. Among them, the Gaussian mixture model stands out as one of the most extensively studied and applied, due to its mathematical tractability and empirical success in sub-population inference, clustering, density estimation, and outlier detection [15, 36].
@@
In the long history of Gaussian mixture models, the identifiability of $\theta$, up to label switching, under model (1.1) is well-known [38, 49]. See also [39] and [28] as well as the recent [2] for identifiability of general finite mixture models. Regarding estimation, a significant line of work analyzes the maximum likelihood estimation (MLE), for density estimation [19], and for estimating the mixing measure, $\rho=\sum_{\ell=1}^{L}\pi_{\ell}\delta_{\{\mu_{\ell},\Sigma\}},$ with $\delta$ being the Dirac measure under the exactly specified (in terms of the number of fitted mixture components) and the over-specified settings [10, 21, 22]. However, since the likelihood function under GMMs is non-concave, computing the MLE is challenging, especially in multivariate settings. To overcome this challenge, three main types of alternative approaches have been developed and studied.
@@
  • Method of Moments (MoM)-based approaches. MoM estimation of $\theta$ has been developed in [29, 23, 5, 20, 34, 26]. Although these methods require minimal separation between Gaussian components, they typically yield slow convergence rates and suboptimal sample complexity with respect to both the dimension $d$ and the number of components $L$. Recently, [45, 16] studied MoM-based estimation of the mixing measure and showed that it achieves the minimax optimal rate under the Wasserstein distance for bounded $L$. While their rates for estimating the mixing measure could be translated into rates for estimating $\theta$ with improved sample complexity, the resulting convergence is still exponentially slow in $L$, even in the univariate case, and corresponding results in the multivariate setting remain largely unknown.
@@
  • Clustering-based approaches. Another substantial body of literature focuses on clustering $X_{1},\ldots,X_{n}$ by predicting their labels $Y_{1},\ldots,Y_{n}$ in model (1.1), a task also known as the community detection problem [12, 37, 41, 27, 3, 1, 30, 11]. Once the labels are known, the parameter $\theta$ can be readily estimated from the labeled data. However, as we detail in Section 4.4, such methods require strong separation conditions between Gaussian components to achieve (nearly) perfect clustering.
@@
  • The Expectation Maximization (EM) algorithm. The EM algorithm is designed in a broader context to approximate the MLE by iteratively maximizing a lower bound of the likelihood function through alternating between the expectation step (E-step) and the maximization step (M-step) [15]. Under the $L$-GMM, both steps have closed-form expressions, leading to fast computation. In contrast to its empirical success, EM is only known to converge to a local stationary point in general, which may be far from the true parameter [44, 36]. Over the past two decades, significant progress has been made in the theoretical understanding of EM’s convergence to the true parameter under the $L$-GMM.
@@
The general positioning of the EM algorithm lies between the first two approaches: compared to MoM-based methods, EM estimation, when suitably initialized, typically requires more separation but yields faster and, oftentimes minimax optimal, rates of convergence; whereas compared to clustering-based methods, EM requires weaker separation condition to achieve optimal convergence rates. The primary goal of this work is to conduct a thorough theoretical analysis of the EM algorithm to formally demonstrate its adaptivity under the $L$-GMM in the presence of mild separation conditions. As explained below, existing analyses of EM are largely limited either to the case $L=2$ or to settings where both the mixing probabilities $\pi$ and the covariance matrix $\Sigma$ are known (see the summary in Table 1). Analyzing EM for mixtures with $L\geq 3$ presents significant challenges, as the likelihood landscape deteriorates with increasing $L$. Furthermore, allowing both $\pi$ and $\Sigma$ to be unknown requires not only estimating these quantities but also adjusting the estimation of the mean vectors. This work is the first to establish both the convergence of EM to the true parameters and its minimax optimality under the $L$-GMM with $L\geq 2$ and both $\pi$ and $\Sigma$ unknown.
@@ 1.1 Related literature
In this section, we review existing theoretical results on the convergence of the EM algorithm under Gaussian mixture models. The early work by [48] establishes a connection between EM iterates and the gradient of the likelihood function under general GMMs. Later, [13] analyzed a two-step EM procedure equipped with a specially designed initialization scheme, under GMMs with $L\geq 2$ components and isotropic covariance matrices. Under a strong separation condition between Gaussian components, the two-step EM estimator is proved to converge to within a small neighborhood of the true parameter, although the size of this neighborhood may not shrink to zero as the sample size increases.
@@
Recently, there has been renewed progress in understanding the convergence of EM to the true parameter under the $L$-GMM. In [4], the authors develop a general framework for analyzing EM and apply it to the symmetric, isotropic 2-GMM with $\pi_{1}=\pi_{2}=1/2$, $\mu_{1}=-\mu_{2}=\mu$, and $\Sigma=\mathbf{I}_{d}$. They prove that the EM iterates $\widehat{\mu}^{(t)}$ converge to the true parameter $\mu$ in $\ell_{2}$-norm at the rate $\|\mu\|_{2}^{3}\sqrt{d/n}$, provided that the initialization $\widehat{\mu}^{(0)}$ satisfies $\|\widehat{\mu}^{(0)}-\mu\|_{2}\leq\|\mu\|_{2}/4$, and the separation between Gaussian components, $\|\mu\|_{2}$, exceeds a sufficiently large constant. Further improvements in this symmetric, isotropic $2$-GMM are made in [47] and [14], where the population level EM is shown to converge to the true parameter as long as the separation $\|\mu\|_{2}>0$ and the initialization $\widehat{\mu}^{(0)}$ is not orthogonal to $\mu$. In other words, the population level likelihood function in this special case has no spurious local maxima, so that any random initialization suffices for the sample level EM algorithm to converge to the ground truth as $n\to\infty$. More recently, still under the setting $\mu_{1}=-\mu_{2}=\mu$ and $\Sigma=\mathbf{I}_{d}$, it has been shown that the EM iterates with simple random initialization converge to the MLE and achieve adaptive minimax optimality for $\|\mu\|_{2}=\mathcal{O}(1)$, with equal mixing weights $\pi_{1}=\pi_{2}=1/2$ [46] and for unequal but known weights [43]. Similar guarantees are also obtained for the case where $L=2$ is specified but the true model consists of a single Gaussian component [17]. These strong global convergence results, however, rely on the assumptions of known $\pi$ and $\Sigma=\mathbf{I}_{d}$. When these conditions are not met, EM must also estimate both $\pi$ and $\Sigma$, and even for $L=2$, it remains unclear whether global convergence guarantees are possible. For the 2-GMM with both unknown $\pi$ and unknown $\Sigma$, [9] established convergence of the EM algorithm to the true parameter, provided that a suitable initialization condition is met and the separation between Gaussian components is sufficiently large but remains bounded.
@@
Despite of these works on the $2$-GMM, they can not extend to the $L$-GMM with $L\geq 3$ from several aspects. First, the landscape of the likelihood function for $L\geq 3$ differs drastically from the two-component case. As shown in [25], the population-level likelihood function under the $L$-GMM with $L\geq 3$ can exhibit poor local maxima even in the simplest case of equally weighted mixtures of well-separated, spherical Gaussians. Moreover, the EM algorithm, when initialized randomly from the data, converges with high probability to such bad local maxima. It is therefore crucial to establish, even at the population level, conditions under which EM converges to the true parameter. Specifically, one must characterize: (1) the region in which EM contracts; (2) the type of initialization that ensures the iterates remain within this region; and (3) how the required separation between Gaussian components depends on $L$, the minimal mixing proportion $\pi_{\min}=\min_{\ell}\pi_{\ell}$ as well as the dimension $d$. Second, these issues must also be addressed at the sample level. Third, once we allow the aforementioned quantities to diverge as $n\to\infty$, both the minimax optimal rate for estimating $\theta$, and whether EM provably achieves this rate, remain open problems.
The only existing works that study the convergence of EM to the true parameter under the $L$-GMM with $L\geq 3$ is Yan et al. [50] for a gradient-based EM and Zhao et al. [51] for the standard EM, both of which assume known $\pi$ and $\Sigma=\mathbf{I}_{d}$. Their results (see Yan et al. [50, Theorems 1 & 4]) require the Euclidean distances between the mean vectors to satisfy
$\min_{k\neq k^{\prime}}\|\mu_{k}-\mu_{k^{\prime}}\|_{2}^{2}~\gg~(L\wedge d)\left\{\log\Bigl{(}\max_{k,k^{\prime}}\|\mu_{k}-\mu_{k^{\prime}}\|_{2}^{2}+L\wedge d\Bigr{)}+\log(1/\pi_{\min})\right\}.$ (1.2)
@@
Given an initialization within a neighborhood of the true parameter, the authors also derive rates of convergence for their estimators. However, our results reveal that the separation condition required for EM to converge can be substantially weaker, even when both $\pi$ and $\Sigma$ are unknown and must be estimated by the algorithm. Moreover, the convergence rates established in [50, 51] are not minimax optimal and can be improved in several respects. Table 1 summarizes a technical comparison with existing works for $L\geq 2$, while a full table including existing works for $L=2$ is deferred to Table 3 in Appendix A.
Table 1: Comparison with existing theoretical results of the EM algorithm and its variant under the $L$-GMM with $L\geq 2$. The notation $\widetilde{\mathcal{O}}$ hides multiplicative logarithmic factors.
@@
Parameter settings Separation $\Delta_{\min}$ Initialization Convergence Rates Optimality
@@
[13] $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\Omega\Bigl{(}\sqrt{d\log\pi_{\min}^{-1}}+\log\pi_{\min}^{-1}\Bigr{)}$ A particular initialization $d(\widehat{\pi},\pi)+d(\widehat{M},M)=\mathcal{O}\left(\sqrt{d\over n\pi_{\min}}+{\exp(-c\sqrt{d})\over\pi_{\min}}\right)$ $\times$
@@
[50] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl{(}(d\wedge L)\log{\Delta_{\max}\over\pi_{\min}}\Bigr{)}$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)=\widetilde{\mathcal{O}}\left({\pi_{\max}\over\pi_{\min}}\sqrt{\Delta_{\max}}(\sqrt{d}+L^{3}\Delta_{\max})\sqrt{d\over n}\right)$ $\times$
@@
[51] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl{(}(d\wedge L)\log{\Delta_{\max}\over\pi_{\min}}\Bigr{)}$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)=\widetilde{\mathcal{O}}\left({\sqrt{\Delta_{\max}}\over\pi_{\min}}\sqrt{dL\over n}\right)$ $\times$
@@
This work Full generality $\Omega\Bigl{(}\log{\Delta_{\max}\over\pi_{\min}}\Bigr{)}$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{\pi},\pi)+d(\widehat{M},M)=\widetilde{\mathcal{O}}\Bigl{(}\sqrt{d\over n\pi_{\min}}\Bigr{)}$ $\checkmark$
$d(\pi^{(0)},\pi)<1,~d(\Sigma^{(0)},\Sigma)\leq c$,
or via preliminary labels
$d(\widehat{\Sigma},\Sigma)=\widetilde{\mathcal{O}}\Bigl{(}\sqrt{d\over n}\Bigr{)}$

### v2

@@ Abstract
Gaussian mixture models (GMMs) are fundamental statistical tools for modeling heterogeneous data. Due to the nonconcavity of the likelihood function, the Expectation-Maximization (EM) algorithm is widely used for parameter estimation of each Gaussian component. Existing analyses of the EM algorithm’s convergence to the true parameter focus on either the two-component case or multi-component settings with known mixing probabilities and isotropic covariance matrices.
In this work, we study the convergence of the EM algorithm for multi-component GMMs in full generality. The population-level EM is shown to converge to the true parameter when the smallest separation among all pairs of Gaussian components exceeds a logarithmic factor of the largest separation and the reciprocal of the minimal mixing probabilities. At the sample level, the EM algorithm is shown to be minimax rate-optimal, up to a logarithmic factor. We develop two distinct novel analytical approaches, each tailored to a different regime of separation, reflecting two complementary perspectives on the use of EM. As a byproduct of our analysis, we show that the EM algorithm, when used for community detection, also achieves the minimax optimal rate of misclustering error under milder separation conditions than spectral clustering and Lloyd’s algorithm, an interesting result in its own right. Our analysis allows the number of components, the minimal mixing probabilities, the separation between Gaussian components and the dimension to grow with the sample size. Simulation studies corroborate our theoretical findings.
Keywords: Community detection; Clustering; EM algorithm; Gaussian mixture models; Minimax optimal estimation.
@@ 1 Introduction
Finite mixture models are fundamental statistical tools for modeling heterogeneous data [lindsay1995mixture, mclachlan1988mixture] and have found wide-ranging applications in fields such as biology, economics, image processing, and natural language processing, to name just a few. These models assume that the observed data arise from a mixture of several latent sub-populations, each governed by its own distribution [mclachlan2000finite, titterington1985statistical]. Among them, the Gaussian mixture model stands out as one of the most extensively studied and applied, due to its mathematical tractability and empirical success in sub-population inference, clustering, density estimation, and outlier detection [dempster1977maximum, redner1984mixture].
@@
In the long history of Gaussian mixture models, the identifiability of $\theta$, up to label switching, under model (1.1) is well-known [teicher1963identifiability, yakowitz1968identifiability]. See also [titterington1985statistical] and [lindsay1995mixture], as well as the recent [allman2009identifiability] for general finite mixture models. Regarding estimation, a significant line of work analyzes the maximum likelihood estimation (MLE) for density estimation [ghosal2001entropies] and for estimating the mixing measure, $\rho=\sum_{\ell=1}^{L}\pi_{\ell}\delta_{\{\mu_{\ell},\Sigma\}},$ with $\delta$ being the Dirac measure under the exactly specified (in terms of the number of fitted mixture components) and over-specified settings [chen1995optimal, ho2016convergence, ho2016strong]. However, since the likelihood function under GMMs is non-concave, computing the MLE is challenging, especially in multivariate settings. To overcome this challenge, three main types of alternative approaches have been developed and studied.
@@
  • Method of Moments (MoM)-based approaches. MoM estimation of $\theta$ has been developed in [LindsayBasak, hsu2013learning, belkin2010polynomial, hardt2015tight, moitra2010settling, kalai2010efficiently]. Although these methods require minimal separation between Gaussian components, they typically yield slow convergence rates and suboptimal sample complexity with respect to both the dimension $d$ and the number of components $L$. Recently, [WuYang2020, doss2023optimal] studied MoM-based estimation of the mixing measure and showed that it achieves the minimax optimal rate under the Wasserstein distance for bounded $L$. While their rates for estimating the mixing measure could be translated into rates for estimating $\theta$ with improved sample complexity, the resulting convergence is still exponentially slow in $L$, even in the univariate case, and corresponding results in the multivariate setting remain largely unknown.
@@
  • Clustering-based approaches. Another substantial body of literature focuses on clustering $X_{1},\ldots,X_{n}$ by predicting their labels $Y_{1},\ldots,Y_{n}$ in model (1.1), a task also known as the community detection problem [dasgupta1999learning, sanjeev2001learning, vempala2002spectral, kannan2008spectral, awasthi2012improved, achlioptas2005spectral, lu2016statistical, chen2024achieving]. Once the labels are known, the parameter $\theta$ can be readily estimated from the labeled data. However, as we detail in Section 4.4, such methods require strong separation conditions between Gaussian components to achieve (nearly) perfect clustering.
@@
  • The Expectation Maximization (EM) algorithm. The EM algorithm is designed in a broader context to approximate the MLE by iteratively maximizing a lower bound of the likelihood function through alternating between the expectation step (E-step) and the maximization step (M-step) [dempster1977maximum]. Under the $L$-GMM, both steps have closed-form expressions, leading to fast computation. In contrast to its empirical success, EM is only known to converge to a local stationary point in general, which may be far from the true parameter [wu1983convergence, redner1984mixture]. Over the past two decades, significant progress has been made in the theoretical understanding of EM’s convergence to the true parameter under the $L$-GMM.
@@
The general positioning of the EM algorithm lies between the first two approaches: compared to MoM-based methods, EM estimation, when suitably initialized, typically requires more separation but yields faster and, oftentimes minimax optimal, rates of convergence; whereas compared to clustering-based methods, EM requires weaker separation condition to achieve optimal convergence rates. The primary goal of this work is to conduct a thorough theoretical analysis of the EM algorithm to formally demonstrate its adaptivity under the $L$-GMM in the presence of mild separation conditions. As explained below, existing analyses of EM are largely limited either to the case $L=2$ or to settings where the mixing probabilities $\pi$ are known and the covariance matrix $\Sigma$ are isotropic (see the summary in Table 1). Analyzing EM for mixtures with $L\geq 3$ presents significant challenges, as the likelihood landscape deteriorates with increasing $L$. Furthermore, allowing $\pi$ and $\Sigma$ to be unknown requires not only estimating these quantities but also adjusting the estimation of the mean vectors, as the E-steps (see Section 2) depend on their estimates as well. This work is the first to establish both the convergence of EM to the true parameter and its minimax optimality under the $L$-GMM with $L\geq 2$ and both $\pi$ and $\Sigma$ unknown.
@@ 1.1 Related literature
In this section, we review existing theoretical results on the convergence of the EM algorithm under Gaussian mixture models. The early work by [xu1996convergence] establishes a connection between EM iterates and the gradient of the likelihood function under general GMMs. Later, [dasgupta2007probabilistic] analyzed a two-step EM procedure equipped with a specially designed initialization scheme, under GMMs with $L\geq 2$ components and isotropic covariance matrices. Under a strong separation condition between Gaussian components, the two-step EM estimator is proved to converge to within a small neighborhood of the true parameter, although the size of this neighborhood may not shrink to zero as the sample size increases.
@@
Recently, there has been renewed progress in understanding the convergence of EM to the true parameter under the $L$-GMM. In [EM2017], the authors develop a general framework for analyzing EM and apply it to the symmetric, isotropic 2-GMM with $\pi_{1}=\pi_{2}=1/2$, $\mu_{1}=-\mu_{2}=\mu$, and $\Sigma=\mathbf{I}_{d}$. They prove that the EM iterates $\widehat{\mu}^{(t)}$ converge to the true parameter $\mu$ in $\ell_{2}$-norm at the rate $\|\mu\|_{2}^{3}\sqrt{d/n}$, provided that the initialization $\widehat{\mu}^{(0)}$ satisfies $\|\widehat{\mu}^{(0)}-\mu\|_{2}\leq\|\mu\|_{2}/4$, and the separation between Gaussian components, $\|\mu\|_{2}$, exceeds a sufficiently large constant. Further improvements in this symmetric, isotropic $2$-GMM are made in [Xu_Hsu_Maleki] and [daskalakis2017ten], where the population level EM is shown to converge to the true parameter as long as the separation $\|\mu\|_{2}>0$ and the initialization $\widehat{\mu}^{(0)}$ is not orthogonal to $\mu$. In other words, the population level likelihood function in this special case has no spurious local maxima, so that any random initialization suffices for the sample level EM algorithm to converge to the ground truth as $n\to\infty$. More recently, still under the setting $\mu_{1}=-\mu_{2}=\mu$ and $\Sigma=\mathbf{I}_{d}$, it has been shown that the EM iterates with simple random initialization converge to the MLE and achieve adaptive minimax optimality for $\|\mu\|_{2}=\mathcal{O}(1)$, with equal mixing weights $\pi_{1}=\pi_{2}=1/2$ [wu2021randomly] and for unequal but known weights [weinberger2022algorithm]. Similar guarantees are also obtained for the case where $L=2$ is specified but the true model consists of a single Gaussian component [dwivedi2020singularity]. These strong global convergence results, however, rely on the assumptions of known $\pi$ and $\Sigma=\mathbf{I}_{d}$. When these conditions are not met, EM must also estimate both $\pi$ and $\Sigma$, and even for $L=2$, it remains unclear whether global convergence guarantees are possible. For the 2-GMM with both unknown $\pi$ and unknown $\Sigma$, [cai2019chime] established convergence of the EM algorithm to the true parameter, provided that a suitable initialization condition is met and the separation between Gaussian components is sufficiently large but remains bounded.
@@
Despite of these works on the $2$-GMM, they can not extend to the $L$-GMM with $L\geq 3$ from several aspects. First, the landscape of the likelihood function for $L\geq 3$ differs drastically from the two-component case. As shown in [jin2016local], the population-level likelihood function under the $L$-GMM with $L\geq 3$ can exhibit poor local maxima even in the simplest case of equally weighted mixtures of well-separated, spherical Gaussians. Moreover, the EM algorithm, when initialized randomly from the data, converges with high probability to such bad local maxima. It is therefore crucial to establish, even at the population level, conditions under which EM converges to the true parameter. Specifically, one must characterize: (1) the region in which EM contracts; (2) the type of initialization that ensures the iterates remain within this region; and (3) how the required separation between Gaussian components depends on $L$, the minimal mixing proportion $\pi_{\min}=\min_{\ell}\pi_{\ell}$ as well as the dimension $d$. Second, these issues must also be addressed at the sample level. Third, once we allow the aforementioned quantities to diverge as $n\to\infty$, both the minimax optimal rate for estimating $\theta$, and whether EM provably achieves this rate, remain open problems.
@@
For such reasons, much fewer results exist on the convergence of the EM algorithm for $L\geq 3$. When $\pi$ is known and $\Sigma=\mathbf{I}_{d}$, yan2017convergence derive the convergence rate for a gradient-based EM algorithm, while Zhao2020 analyze the standard EM. Their results require that $\min_{k\neq\ell}\|\mu_{k}-\mu_{\ell}\|_{2}$ satisfy a certain separation condition and that the initialization lies within a neighborhood of the true parameter. This separation condition is subsequently relaxed in [kwon2020algorithm, regev2017learning], and the rate analyses are refined by [segol2021improved]. However, none of these works apply when both $\pi$ and $\Sigma$ are unknown. Moreover, the convergence rates established in the aforementioned studies are not minimax-optimal, even under known $\pi$ and $\Sigma$, and can be improved in several respects. Table 1 summarizes a technical comparison with existing results for $L\geq 2$, while a full table including the $L=2$ case is deferred to Table 3 in Appendix A.
[…]
@@
Settings Separation $\Delta_{\min}\gtrsim$ Initialization Convergence Rates Optimality
@@
[dasgupta2007probabilistic] $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\sqrt{d\log(\pi_{\min}^{-1})}+\log(\pi_{\min}^{-1})$ A particular initialization $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lesssim\sqrt{d/(n\pi_{\min})}+\exp(-c\sqrt{d})/\pi_{\min}$ $\times$
[kwon2020algorithm] $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ A sample-splitting version of the EM $\checkmark$
@@
[yan2017convergence] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $(d\wedge L)\log(\Delta_{\max}/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(\pi_{\max}/\pi_{\min})\sqrt{\Delta_{\max}}(\sqrt{d}+L^{3}\Delta_{\max})\sqrt{d/n}$ $\times$
@@
[Zhao2020] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $(d\wedge L)\log(\Delta_{\max}/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(\sqrt{\Delta_{\max}}/\pi_{\min})\sqrt{dL/n}$ $\times$
[segol2021improved] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(1/\pi_{\min})\sqrt{dL/n}$ $\times$
@@
This work Full generality $\log(\Delta_{\max}/\pi_{\min})$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lessapprox\sqrt{d/(n\pi_{\min})}$ $\checkmark$
$d(\Sigma^{(0)},\Sigma)\leq c$, or via preliminary labels $d(\widehat{\Sigma},\Sigma)\lessapprox\sqrt{d/n}$